# Growth and collapse of subsystem complexity under random unitary circuits

### v1

@@ A.2 Purity of a finite interval on an infinite line
where $C$ is a configuration of binary variables (0 or $1$), one for each of the random gates in the circuit and sites located at the top in a spacetime diagram, subject to a boundary condition that all site binary variables inside the interval are 0 (up) and all others $1$ (down) and where the weight $|\partial C|$ is the length of domain wall, calculated by the rule shown in Fig.˜2.
Figure 2: Two domains walls propagate from the top to bottom and never go up. They may or may not merge in the bulk. Each have weight 77 in the figure, giving total weight $14$.
There are domain configurations with a single-component domain wall, as well as those whose end points lie at the bottom. We always think of domain walls propagating from the top to the bottom. The weight of a configuration is determined by the time coordinate of the merger point. If the merger happens at the slice of time coordinate $t>0$, the length of the domain wall is 2(Tt)<2T2(T-t)<2T. If two domain walls reach $t=0$, then they may or may not meet, but in any case the total length of the domain walls is $2T$.
@@
Figure 3: Coordinate system for counting the domain walls. The points $C$ and $D$ are present only if the two domain walls merge. The meeting point is neither $C$ nor $D$; these are points immediately before the merger.
@@
This may be proved by tedious binomial coefficient calculation, but a combinatorial proof is possible as follows. For any even integer $z\geq 0$, let ${\mathcal{N}}(z)$ be the set of all paths that start from $(\frac{z}{2},0)$ and $(0,\frac{z}{2})$ and merge at some point by moving up and right. The number of elements of ${\mathcal{N}}(z)$ with total length $x+y+z$ is $N(x,y;z)$. Take an element $\gamma\in{\mathcal{N}}(\ell+2)$ merging at $(x^{\prime},y^{\prime})$ with $x^{\prime}+y^{\prime}=f+\ell+2$. Let $g=x+y$ be the least such that (x1,y)(x-1,y) and (x,y1)(x,y-1) are on $\gamma$. Since the subpaths of $\gamma$ meet at $(x^{\prime},y^{\prime})$ for the first time, we must have $g\leq f+\ell<x^{\prime}+y^{\prime}$. Cut at $x+y=g$ line, the path $\gamma$ splits into three components. The top-right component defines an element of ${\mathcal{N}}(2)$. If we shift the upper left component to the right by $1$ and shift the lower right component up by $1$, then these shifted components merge at $(x,y)$ but nowhere else since $g$ was chosen to be the least. This shift-merged path defines an element of ${\mathcal{N}}(\ell)$. Conversely, given $\gamma^{\prime}\in{\mathcal{N}}(\ell)$ and $\gamma^{\prime\prime}\in{\mathcal{N}}(2)$ where the sum of total lengths is $2(f+\ell+2)$, we can break the tip (the merger point) of $\gamma^{\prime}$, shift the separated components left and down, cap them with $\gamma^{\prime\prime}$, and obtain an element of ${\mathcal{N}}(\ell+2)$. This proves Eq.˜A.10 and in turn the second claim.
@@
${\tfrac{1}{2}}-{\tfrac{1}{2}}(1-4\eta^{2})^{1/2}=\int_{0}^{\eta^{2}}(1-4u)^{-{\tfrac{1}{2}}}{\mathrm{d}}u$ $=\sum_{k=0}^{\infty}\eta^{2(k+1)}\frac{1}{k+1}\binom{2k}{k}$
$=\eta^{2}+\sum_{k=1}^{\infty}\eta^{2(k+1)}\frac{1}{k}\binom{2k}{k+1}$ (A.12)
$=\eta^{2}+\sum_{k=1}^{\infty}\eta^{2(k+1)}\frac{1}{k}\binom{2k}{k-1}\,$.
@@
Define
$N(z;\ell)$ $=\sum_{x,y\,:\,x+y=z}N(x,y;\ell)\,.$ (A.15)
Note that $N(z;\ell)=0$ if $z<\frac{\ell}{2}$ and $N(\tfrac{\ell}{2};\ell)=1$. Next, we examine part of domain wall restricted to the region where $x+y\leq T+\frac{\ell}{2}$. Define
${\mathcal{J}}(T;\ell)$ $=\left\{\gamma|_{x+y\leq T+\frac{\ell}{2}}\,\middle|\,\gamma\in{\mathcal{N}}(\ell)\right\}\,$, (A.16)
$J(T;\ell)$ $=\#{\mathcal{J}}(T;\ell)\leq 4^{T}\,.$
If $T<\frac{\ell}{2}$, then $J(T;\ell)=4^{T}$. The purity of the interval at time $T$ is then
@@
(Upper bound) The sum over $z$ in Eq.˜A.17 is of nonnegative numbers and converges to $q^{-\ell}$ as $T\to\infty$ by Lemma˜A.2. Therefore, $P(T;\ell)-q^{-\ell}\leq J(T;\ell)\eta^{2T}\leq 4^{T}\eta^{2T}$.
(Lower bound) Let ${\mathcal{J}}(r;T;\ell)\subset{\mathcal{J}}(T;\ell)$ be the set of all configurations such that the two end points on the line $x+y=T$ of the domain wall are separated by $\ell_{1}$-distance $r$. The distance $r$ is even and $0\leq r\leq 2T+\ell$. Let $J(r;T;\ell)=\#{\mathcal{J}}(r;T;\ell)$. If we apply Lemma˜A.2 to the two end points of $\gamma\in{\mathcal{J}}(r;T;\ell)$, we see that
@@
Since $\bigcup_{r}{\mathcal{J}}(r;T;\ell)={\mathcal{J}}(T;\ell)$, we see $J(T;\ell)\eta^{2T}=\sum_{r}J(r;T;\ell)\eta^{2T}$. Therefore,
$P(T;\ell)-q^{-\ell}=\sum_{r=0}^{\ell+2T}J(r;T;\ell)(1-q^{-r})\eta^{2T}\geq J(\ell;T;\ell)(1-q^{-2})\eta^{2T}\,.$ (A.19)
@@ Appendix B Tight norm conversions
Let us show that the inquality for two normalized density matrices $\alpha,\beta$
$\lVert\alpha-\beta\rVert_{1}\leq 2\sqrt{\min(\operatorname{{rank}}\alpha,\operatorname{{rank}}\beta)}\lVert\alpha-\beta\rVert_{2}$ (B.1)
cannot be improved even if we constrain Renyi-($m\geq 1$) entropies. Let $a=\operatorname*{{diag}}(a_{1}\geq a_{2}\geq\cdots a_{r}>0=\cdots=0)$ be a diagonal density matrix of rank $r$ and some large dimension $d$. Let $0<\epsilon<a_{r}$, and define $b=\operatorname*{{diag}}(b_{1}\geq b_{2}\geq\cdots b_{d})$ by
@@
so (B.1) is tight. On the other hand,
@@ Appendix C Shallow unitary designs as a model of chaos
A unitary $k$-design (a probabilistic ensemble of unitaries that is close to the Haar random one) is sometimes taken as a model of chaotic dynamics [7]. Here we explain significant differences between shallow unitary designs [33, 39, 17] and the brickwork random circuits in the phenomenology of subregion complexity. We point out that the complexity of the reduced density matrix of an interval of length $\ell\ll L$ at depth $T\gg\ell$, is of order $T\ell$ under the $\epsilon$-approximate unitary $k$-designs constructed in [39]. Importantly, $T$ here can be exponentially large in $\ell$. This contrasts sharply to the brickwork random circuit, under which the complexity at such late times is zero since the state is the maximally mixed state up to a trace-distance error that is exponentially small in $\ell$.
The circuit $U$ constructed in [39] can be thought of as a 1d circuit of supergates with depth $2$ over $\operatorname{poly}(Lk/\epsilon)$-dimensional superqudits. (A superqudit consists of $\mathcal{O}(\log_{q}(Lk/\epsilon))$ $q$-dimensional qudits.) It follows that the rank of the reduced density matrix $\sigma$ on any interval of length $\ell$ after applying $U$ to a product state is only $r=\operatorname{poly}(Lk/\epsilon)$.
@@
To lower-bound the complexity of typical $\sigma$, we follow the same line of reasoning as in Section˜2. In particular, Eq.˜2.6 becomes
$\operatorname{{\mathbb{E}}}_{\sigma}F(\rho,\sigma)^{2k}\leq(1+\epsilon)r^{k}q^{-k\ell}k!\leq(1+\epsilon)\left(\frac{\operatorname{poly}(Lk/\epsilon)}{q^{\ell}}\right)^{k}.$ (C.1)
Each supergate can be chosen to be the random brickwork circuit on a superqudit, and the depth of a supergate measured in terms of ${\mathsf{U}}(q^{2})$ gates is $T/2$. The corresponding parameter $k$ is proportional to $T$. Setting $\ell=\Theta(L)$, we conclude that typical $\sigma$ has complexity
$\mathfrak{C}(\sigma)=\tilde{\Omega}(\ell T)\text{ for }T\leq q^{\mathcal{O}(\ell)}.$ (C.2)
One may wonder how this is consistent with the purity calculation, which is only quadratic in the density operator:
@@
Since this calculation only involves $(U\otimes U^{\dagger})^{\otimes 2}$, it suffices for $U$ to be an exact unitary $2$-design. However, with $\epsilon$-approximate unitary designs, the error parameter $\epsilon$ becomes very important. Recall that in the defining inequality for the relative error $\epsilon$ in Eq.˜2.1, the observable $M$ is required to be positive semi-definite; however, ${\mathrm{SWAP}}_{A,A^{\prime}}$ has negative eigenvalues. One can decompose $\mathrm{SWAP}$} into positive and negative parts, use Eq.˜2.1, and combine the results, but this gives a completely meaningless result unless $\epsilon$ is exponentially small in $\ell$. In fact, such a bound must not imply that $\rho_{A}$ is close to the maximally mixed state as we have given a lower bound on the complexity using a special approximate unitary design. This is a drastic instance of a sign problem. In the brickwork random circuit, we have used separate calculation (Proposition˜3.1) of purity, which is not related to the unitary $2$-design property.
Note that the random circuits of [33, 39, 17] do not define a simple Markovian random walk on the unitary group. That is, if we have a time sequence of unitaries $U(T)$ of depth $T=1,2,3,\ldots$, constructed by the prescription of those references, then $U(T+1)U(T)^{\dagger}$ is not a quantum circuit of constant depth. Hence, it seems absurd to use these constructions to model time dynamics. If we demand the simple Markovian property, i.e., the distribution of $U(T+1)$ must be the convolution of some distribution of constant depth quantum circuit and the distribution of $U(T)$, then, perhaps, the complexity of any smaller-than-half subsystem would stop growing at time $\Theta(\ell)$.

### v2

@@ A.2 Purity of a finite interval on an infinite line
where $C$ is a configuration of binary variables (0 or $1$), one for each of the random gates in the circuit and sites located at the top in a spacetime diagram, subject to a boundary condition that all site binary variables outside the interval are 0 (up) and all others $1$ (down) and where the weight $|\partial C|$ is the length of domain wall, calculated by the rule shown in Fig.˜2.
Figure 2: Two domains walls must propagate from the top to bottom and must not go up. They may or may not merge in the bulk. Each have weight 77 in the figure, giving total weight $14$.
There are domain configurations with a single-component domain wall or two-component domain wall, in the latter case of which the end points lie at the bottom. We always think of domain walls propagating from the top to the bottom. The weight of a configuration is determined by the time coordinate of the merger point. If the merger happens at the slice of time coordinate $t>0$, the length of the domain wall is 2(Tt)<2T2(T-t)<2T. If two domain walls reach $t=0$, then they may or may not meet, but in any case the total length of the domain walls is $2T$.
@@
Figure 3: Coordinate system for counting the domain walls. The points $C$ and $D$ are present only if the two domain walls merge. The meeting point is neither $C$ nor $D$; these points are immediately before the merger and are used to evaluate $N(x,y;\ell)$.
@@
This may be proved by tedious binomial coefficient calculation, but a combinatorial proof is possible as follows. For any even integer $z\geq 0$, let ${\mathcal{N}}(z)$ be the set of all paths that start from $(\frac{z}{2},0)$ and $(0,\frac{z}{2})$ and merge at some point by moving up and right. Take an element $\gamma\in{\mathcal{N}}(\ell+2)$ merging at $(x^{\prime},y^{\prime})$ with $x^{\prime}+y^{\prime}=f+\ell+2$. Let $g=x+y$ be the least such that (x1,y)(x-1,y) and (x,y1)(x,y-1) are on $\gamma$. Since the subpaths of $\gamma$ meet at $(x^{\prime},y^{\prime})$ for the first time, we must have $g\leq f+\ell<x^{\prime}+y^{\prime}$. Cut at $x+y=g$ line, the path $\gamma$ splits into three components. The top-right component defines an element of ${\mathcal{N}}(2)$. If we shift the upper left component to the right by $1$ and shift the lower right component up by $1$, then these shifted components merge at $(x,y)$ but nowhere else since $g$ was chosen to be the least. This shift-merged path defines an element of ${\mathcal{N}}(\ell)$. Conversely, given $\gamma^{\prime}\in{\mathcal{N}}(\ell)$ and $\gamma^{\prime\prime}\in{\mathcal{N}}(2)$ where the sum of total lengths is $2(f+\ell+2)$, we can break the tip (the merger point) of $\gamma^{\prime}$, shift the separated components left and down, cap them with $\gamma^{\prime\prime}$, and obtain an element of ${\mathcal{N}}(\ell+2)$. This proves Eq.˜A.12 and in turn the second claim.
@@
${\tfrac{1}{2}}-{\tfrac{1}{2}}(1-4\eta^{2})^{1/2}=\int_{0}^{\eta^{2}}(1-4u)^{-1/2}{\mathrm{d}}u$ $=\sum_{k=0}^{\infty}\eta^{2(k+1)}\frac{1}{k+1}\binom{2k}{k}$
$=\eta^{2}+\sum_{k=1}^{\infty}\eta^{2(k+1)}\frac{1}{k}\binom{2k}{k-1}\,$. (A.14)
@@
Next, we examine part of domain wall restricted to the region where $x+y\leq T+\frac{\ell}{2}$. Define
$N(z;\ell)$ $=\sum_{x,y\,:\,x+y=z}N(x,y;\ell)=\begin{cases}0&\text{if }z<\ell\\ 1&\text{if }z=\ell\end{cases}\,$,
[…]
$J(T;\ell)$ $=\#{\mathcal{J}}(T;\ell)\quad\text{ which is }\begin{cases}\leq 4^{T}&\text{always}\\ =4^{T}&\text{if }T<\frac{\ell}{2}\end{cases}\,.$
Since the domain walls either remain disjoint or merge, the purity of the interval at time $T$ is then
@@
(Upper bound) The sum over $z$ in Eq.˜A.18 is of nonnegative numbers and converges to $q^{-\ell}$ as $T\to\infty$ by Lemma˜A.9. Therefore, $P(T;\ell)-q^{-\ell}\leq J(T;\ell)\eta^{2T}\leq 4^{T}\eta^{2T}$.
(Lower bound) Let ${\mathcal{J}}(r;T;\ell)\subset{\mathcal{J}}(T;\ell)$ be the set of all configurations such that the two end points on the line $x+y=T$ of the domain wall are separated by $\ell_{1}$-distance $r$. The distance $r$ is even and $0\leq r\leq 2T+\ell$. Let $J(r;T;\ell)=\#{\mathcal{J}}(r;T;\ell)$. If we apply Lemma˜A.9 to the two end points of $\gamma\in{\mathcal{J}}(r;T;\ell)$, we see that
@@
Since ${\mathcal{J}}(T;\ell)=\bigcup_{r}{\mathcal{J}}(r;T;\ell)$, we see $J(T;\ell)\eta^{2T}=\sum_{r}J(r;T;\ell)\eta^{2T}$. Therefore,
$P(T;\ell)-q^{-\ell}=\sum_{r=0}^{\ell+2T}J(r;T;\ell)(1-q^{-r})\eta^{2T}\geq J(\ell;T;\ell)(1-q^{-\ell})\eta^{2T}\,.$ (A.20)
@@ Appendix B Tight norm conversions
Let us show that the inquality for two normalized density matrices $\rho,\sigma$
$\lVert\rho-\sigma\rVert_{1}\leq 2\sqrt{\min(\operatorname{{rank}}\rho,\operatorname{{rank}}\sigma)}\lVert\rho-\sigma\rVert_{2}$ (B.2)
cannot be improved even if we constrain Renyi-($m\geq 1$) entropies. Let $a=\operatorname*{{diag}}(a_{1}\geq a_{2}\geq\cdots\geq a_{r}>0=\cdots=0)$ be a diagonal density matrix of rank $r$ and some large dimension $d$. Let $0<\epsilon<a_{r}$, and define $b=\operatorname*{{diag}}(b_{1}\geq b_{2}\geq\cdots b_{d})$ by
@@
so (B.2) is tight in the limit of large $d$. On the other hand,
@@ Appendix C Shallow unitary designs as a model of chaos
A unitary $k$-design (a probabilistic ensemble of unitaries that is close to the Haar random one) is sometimes taken as a model of chaotic dynamics [7]. Here we explain significant differences between “shallow” unitary designs [33, 39, 17] and the brickwork random circuits in the phenomenology of subregion complexity. We point out that the complexity of the reduced density matrix of an interval of length $\ell\ll L$ (say $\ell=L/100$) at depth $T\gg\ell$, is of order $T$ under the $\epsilon$-approximate unitary $k$-designs constructed in [39]. Importantly, $T$ here can be exponentially large in $\ell$. This contrasts sharply to the brickwork random circuit, under which the complexity at such late times is zero since the state is the maximally mixed state up to a trace-distance error that is exponentially small in $\ell$.
(a) Brickwork circuit in 1+11+1d
(b) Same as (a) but with the gates straddling between the supergates deleted.
Figure 4: With gates along $2\xi$-separated columns removed, we still have an approximate unitary $k$-design where $k$ is proportional to the depth. Starting with a pure product state, the state of an interval under (b) has much smaller entropy than that under (a).
Consider a 1+11+1d circuit $U$ of supergates with depth $2$ over $q^{\xi}$-dimensional superqudits, where every supergate acts on neighboring two superqudits and is a random unitary drawn from an $\epsilon_{0}$-approximate unitary $k$-design. For sufficiently large $\xi$, the circuit $U$ is an approximate unitary $k$-design, where the approximation error $\epsilon$ is bounded above by iteratively applying the following.
Lemma C.1 (Lemma 8 of [39]).
Let $U_{AB}$ and $U_{BC}$ be independent approximate unitary $k$-designs of errors $\epsilon_{AB}$ and $\epsilon_{BC}$ acting on subsystems $AB$ and $BC$, respectively. Assume that $k^{2}d_{B}^{-1}\leq\tfrac{1}{2}$ where $d_{B}$ is the Hilbert space dimension of $B$. Then, the product $U_{AB}U_{BC}$ is an $\epsilon_{ABC}$-approximate unitary $k$-design where
$1+\epsilon_{ABC}\leq(1+\epsilon_{AB})(1+\epsilon_{BC})e^{\frac{k^{2}}{2d_{B}}}\left(1+\frac{k^{2}}{d_{ABC}}\right)\left(1-\frac{k^{2}}{2d_{AB}}\right)^{-1}\left(1-\frac{k^{2}}{2d_{BC}}\right)^{-1}\,.$ (C.2)
Let $x=k^{2}/d_{B}\leq\frac{1}{2}$. For any positive real $z$, we use the notation $z^{\prime}=\ln(1+z)$. Since $d_{AB}\geq d_{B}$ and $d_{BC}\geq d_{B}$, all the $k$-involving factors are at most $1+x$. Taking log of (C.2) we have that
$\epsilon^{\prime}_{ABC}+2x^{\prime}\leq(\epsilon^{\prime}_{AB}+2x^{\prime})+(\epsilon^{\prime}_{BC}+2x^{\prime})\,.$ (C.3)
Applying (C.3) iteratively, once for each supergate, we see that the full circuit $U$ of $m$ supergates is an $\epsilon$-approximate unitary $k$-design where $\epsilon^{\prime}\leq m\epsilon^{\prime}_{0}+2mx^{\prime}$ if $k^{2}q^{-\xi}\leq\frac{1}{2}$. Use $m\leq L/\xi$ and $z^{\prime}\leq z$, and choose $\xi=c\ell=c^{2}L$ with $c=1/100$. Then, the approximation error $\epsilon$ of $U$ as a unitary $k$-design satisfies
$\ln(1+\epsilon)\leq c^{-2}\epsilon_{0}+2c^{-2}k^{2}q^{-c\ell}\,.$ (C.4)
Each random supergate can be chosen to be a brickwork circuit consisting of elements of ${\mathsf{U}}(q^{2})$ in $T/2$ layers. So, the full circuit $U$ has depth $T$. The result of [13] implies that each supergate is an $\epsilon_{0}$-approximate unitary $k_{0}$-design for some $k_{0}\geq T/\operatorname{poly}(\xi)$ with $\epsilon_{0}=c^{3}$. Setting $k=T/\operatorname{poly}(\xi)$, we have $\epsilon^{\prime}\leq c+2q^{-c\ell}T^{2}/\operatorname{poly}(c\ell)$, which is less than $0.1$ if $T<T_{max}=\tilde{\mathcal{O}}(q^{c\ell/2})$. [^10] In [39], the overlap size $\xi$ was taken to be $\mathcal{O}(\log_{q}(Lk/\epsilon))$ to minimize $T$ given $k$; with this choice of $\xi$ for a given $k$, the overall depth $T$ can be chosen as $\tilde{\mathcal{O}}(k\log(Lk/\epsilon))$. This is why it is a “shallow” unitary design. Note that the overall circuit $U$ is nothing but a brickwork circuit with some ${\mathsf{U}}(q^{2})$ gates deleted along time-like columns separated by $2\xi$. See Fig.˜4.
Due to the deletion, the rank $r$ of the reduced density matrix $\sigma$ on any interval of length $\ell$ after applying $U$ to a product pure state is bounded as
$r\leq q^{4\xi}\ll q^{\ell}\,.$ (C.5)
@@
(Note that $\lVert\sigma-q^{-\ell}{\mathbf{1}}_{q^{\ell}}\rVert_{1}>\frac{1}{2}$ since $r\leq\frac{1}{2}q^{\ell}$.) Then, we follow the same line of reasoning as in Section˜2 for a complexity lower-bound. In particular, Eq.˜2.7 becomes
$\operatorname{{\mathbb{E}}}_{\sigma}F(\rho,\sigma)^{2k}\leq(1+\epsilon)r^{k}q^{-k\ell}k!\leq 2\left(kq^{-(1-4c)\ell}\right)^{k}.$ (C.6)
Setting $k=T/\operatorname{poly}(c\ell)$, we conclude that typical $\sigma$ has complexity
$\mathfrak{C}(\sigma)=\Omega(T/\operatorname{poly}(c\ell,q))\text{ for }T\leq T_{max}=q^{\mathcal{O}(\ell)}.$ (C.7)
So the rank of $\sigma$ is small, and its complexity can be large. These results contrast with the case where $U$ is Haar random — in that case $\sigma$ would be close to maximally mixed, with low complexity. How are these opposite conclusions compatible with the fact that the supergate circuit generates an approximate $k$-design? To temporarily sharpen the puzzle, note that for the Haar random case, these facts can be deduced by computing the purity, which is only quadratic in the density operator:
@@
Since this calculation only involves $(U\otimes U^{\dagger})^{\otimes 2}$, it suffices for $U$ to be an exact unitary ($k=2$)-design.
However, with $\epsilon$-approximate unitary designs, the error parameter $\epsilon$ becomes very important. Recall that in the defining inequality for the relative error $\epsilon$ in Eq.˜2.1, the observable $M$ is required to be positive semi-definite, but ${\mathrm{SWAP}}_{A,A^{\prime}}$ has negative eigenvalues. One can decompose $\mathrm{SWAP}$} into positive and negative parts, use Eq.˜2.1, and combine the results, but this gives a completely meaningless result unless $\epsilon$ is exponentially small in $\ell$. (In view of (C.7), such a small $\epsilon$ will require $c=\xi/\ell$ to be “big.”) In fact, a bound on trace distance with a constant error $\epsilon$ for an approximate unitary design must not imply that $\rho_{A}$ is close to the maximally mixed state since we have given a lower bound on the complexity using a special approximate unitary design. This is a drastic instance of a sign problem. In the brickwork random circuit, we have used separate calculation (Proposition˜3.1) of purity, which is not related to the unitary $2$-design property.
Note that the random circuits of [33, 39, 17] do not define a simple Markovian random walk on the unitary group. That is, if we have a time sequence of unitaries $U(T)$ of depth $T=1,2,3,\ldots$, constructed by the prescription of those references, then $U(T+1)U(T)^{\dagger}$ is not a quantum circuit of constant depth. Hence, it seems absurd to use these constructions to model time dynamics. If we demand the simple Markovian property, i.e., the distribution of $U(T+1)$ must be the convolution of some distribution of constant depth quantum circuits and the distribution of $U(T)$, then, perhaps, the complexity of any smaller-than-half subsystem would stop growing at time $\Theta(\ell)$.